# Self-NPO: Negative Preference Optimization of Diffusion Models by Simply Learning from Itself without Explicit Preference Annotations

### v1

@@ Abstract
Diffusion models have demonstrated remarkable success in various visual generation tasks, including image, video, and 3D content generation. Preference optimization (PO) is a prominent and growing area of research that aims to align these models with human preferences. While existing PO methods primarily concentrate on producing favorable outputs, they often overlook the significance of classifier-free guidance (CFG) in mitigating undesirable results. Diffusion-NPO addresses this gap by introducing negative preference optimization (NPO), training models to generate outputs opposite to human preferences and thereby steering them away from unfavorable outcomes. However, prior NPO approaches, including Diffusion-NPO, rely on costly and fragile procedures for obtaining explicit preference annotations (e.g., manual pairwise labeling or reward model training), limiting their practicality in domains where such data are scarce or difficult to acquire. In this work, we introduce Self-NPO, a Negative Preference Optimization approach that learns exclusively from the model itself, thereby eliminating the need for manual data labeling or reward model training. Moreover, our method is highly efficient and does not require exhaustive data sampling. We demonstrate that Self-NPO integrates seamlessly into widely used diffusion models, including SD1.5, SDXL, and CogVideoX, as well as models already optimized for human preferences, consistently enhancing both their generation quality and alignment with human preferences.
1 Introduction
Over the past few years, diffusion models ho2020denoising ; song2020denoising ; song2020score have made significant strides in various visual generation tasks, including image karras2022elucidating ; rombach2022high ; wang2024rectified ; podell2023sdxl ; diffusionbeatgan ; meng2021sdedit ; teng2024dim ; ma2024storynizor ; li2024ecnet ; sun2024journeydb , video blattmann2023align ; mao2024osv ; shi2024motion ; makeavideo ; chen2024videocrafter2 ; bian2025gs ; NEURIPS2022_f7f47a73 ; wang2024animatelcm ; wang2025your, and 3D generation gao2024cat3d ; poole2022dreamfusion ; li2025connecting ; yan2025consistent ; sun2024genca ; lai2025unleashing. However, diffusion models trained on massive unfiltered data (e.g., image-text pairs) often generate results that do not align well with human preferences. The growing interest in aligning these models for human preference-aligned generation has led to the development of preference optimization (PO) methods. In general, current PO methods can be categorized into four types:
@@ 1 Introduction
Differentiable Reward (DR): In this approach, images generated via iterative denoising are evaluated by a pre-trained reward model, and the diffusion models are subsequently refined through gradient-based backpropagation to align with the reward model xu2024imagereward ; prabhudesai2024video ; zhang2024large ; wu2023hpsv1 ; wu2023human ; wu2024deepreward .
Reinforcement Learning (RL): Here, the denoising process in diffusion models is treated as a Markov decision process (MDP) puterman2014markov ; sutton2018reinforcement . Techniques like PPO and its variants are employed to optimize preferences schulman2017proximal . Images are generated and evaluated dynamically based on reward feedback, with the goal of maximizing the chances of producing images that yield higher rewards.
Direct Preference Optimization (DPO): DPO streamlines training by reframing the reinforcement learning objective into a simulation-free framework rafailov2024direct ; wallace2024diffusion , eliminating the need for separate reward models or online image generation and evaluation. Instead, it relies on fine-tuning with existing datasets of paired preference annotations. Some works yuan2024self ; deng2024enhancing ; gu2024diffusion ; zhang2024seppo construct preference pairs by using high-quality demonstration data as positive examples and self-generated data as negative examples. We have more discussion with these works as they also mention the idea of self-generated data.
Negative Preference Optimization (Diffusion-NPO): Despite previous efforts to make models generate human-aligned images, Diffusion-NPO wang2025diffusionnpo firstly raised an important question: How can a model know to avoid generating poor images if it only knows how to generate good ones without understanding what is bad? Diffusion-NPO introduced the concept of negative preference optimization (NPO), which trains models to generate outputs that oppose human preferences, effectively improving performance by guiding the model away from undesired results though the classifier-free guidance (CFG) ho2022classifier ; karras2024guiding ; Shen_2024_CVPR ; ahn2024self .
@@
However, we argue that a major limitation of existing methods including DR, RL, DPO, and Diffusion-NPO, is their reliance on explicit preference annotations, such as the expensive manual collection of preference data pairs kirstain2023pick ; wu2023hpsv2 and the challenging and fragile process of reward model training xu2024imagereward ; wu2023hpsv2 ; schuhmann2022laion . This heavy dependency on preference annotations makes these approaches less practical, particularly in domains where obtaining or labeling preference data is expensive or infeasible.
@@
In this work, we introduce Self-NPO, a novel approach for negative preference optimization (NPO) that can be done in an explicit preference annotation-free manner. Specifically, our training objective is to apply negative preference optimization to a pre-trained model, making its predictions more aligned with the opposite direction of human preferences. This facilitates unconditional/negative-conditional outputs in CFG, thus reducing the probability of generating results that do not align with human preferences.
@@
    To avoid altering the variance of the epsilon predictions, $\boldsymbol{\epsilon}_{\boldsymbol{\theta}_{neg}}(\mathbf{x}_{t},t,\boldsymbol{c^{\prime}})$ should maintain a certain level of correlation with $\boldsymbol{\epsilon}_{\boldsymbol{\theta}_{pos}}(\mathbf{x}_{t},t,\boldsymbol{c})$. For example, for two completely independent Gaussian noises, the above operation would result in a variance of $12\omega^{2}+2\omega+$ 1.
@@
    Efficient tuning without full diffusion. We do not need to fully execute iterative diffusion generation process to implement self-learning. We propose a novel tuning strategy, which allows us to fine tune diffusion models on the partially excuted iterative diffusion generation results. This avoids the huge generation costs of preparing data for self-learning.
@@
We demonstrate that Self-NPO can be seamlessly integrated with existing models, such as SD1.5 rombach2022high , SDXL podell2023sdxl , and CogVideoX yang2024cogvideox , as well as models previously fine-tuned through preference optimization, consistently improving their alignment with human preferences.
Figure 2: Motivation of Self-NPO. Preference Optimization (PO) optimizes the preference aligned model by training on preference annotations. Negative Preference Optimization (NPO) optimizes the negative preference aligned model by training on switched preference annotations. Self-NPO optimizes the negative preference aligned model by truncated diffusion fine-tuning on truncated self-generated data.
2 Preliminary
@@
Preliminary of CFG. CFG has become a necessary and important technique for improving generation quality and text alignment of diffusion models. For convenience, we focus our discussion on the general formal of diffusion models, i.e., $\boldsymbol{x}_{t}=\alpha_{t}\boldsymbol{x}_{0}+\sigma_{t}\boldsymbol{\epsilon}$ kingma2021variational . Suppose we learn a score estimator from a epsilon prediction neural network $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},\boldsymbol{c},t)$, and we have $\nabla_{\boldsymbol{x}_{t}}\log\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c};t)=-\frac{\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t)}{\sigma_{t}}$. The sample prediction at timestep $t$ of the score estimator is formulated as
@@
$\hat{\boldsymbol{x}}_{0}=\frac{1}{\alpha_{t}}(\boldsymbol{x}_{t}+\sigma^{2}\nabla_{\boldsymbol{x}_{t}}\log\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c};t))\,.$ (2)
@@
Applying the CFG is equivalent to add an additional score term, that is, we replace $\nabla_{\boldsymbol{x}_{t}}\log\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c};t)$ in Eq. 2 with the following term,
$\begin{split}\nabla_{\boldsymbol{x}_{t}}\log\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c};t)+\nabla_{\boldsymbol{x}_{t}}\log\left[\frac{\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c};t)}{\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c}^{\prime};t)}\right]^{\omega}\,,\end{split}$ (3)
@@
where $\omega$ is to control the strength of CFG, $\boldsymbol{c}$ and $\boldsymbol{c}^{\prime}$ are conditional and unconditional/negative-conditional inputs, respectively. It is apparent that the generation will be pushed to high probability region of $\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c};t)$ and relatively low probability region of $\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c}^{\prime};t)$. Write the above equation into the epsilon format, and then we have
$\boldsymbol{\epsilon}^{\omega}_{\boldsymbol{\theta}}=(\omega+1){\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},\boldsymbol{c},t)}-\omega{\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},\boldsymbol{c}^{\prime},t)}\,.$ (4)
@@
Diffusion-NPO wang2025diffusionnpo . Diffusion models use classifier-free guidance (CFG) to improve generation quality by combining conditional and negative-conditional outputs to favor preferred results. Traditional preference optimization aligns outputs with human preferences but overlooks avoiding undesirable outputs. Negative Preference Optimization (NPO), introduced in Diffusion-NPO, trains an additional model to counteract human preferences, reducing undesired outputs via CFG.
@@ 3 Methodology
3.1 NPO via Self-Generated Data
Reinforcement Learning with Human Feedback (RLHF) rafailov2024direct ; wallace2024diffusion ; griffith2013policy ; liu2025improving aims to optimize a conditional distribution $\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})$, where $\boldsymbol{c}\sim\mathcal{D}_{c}$, in order to maximize the expectation value of the associated reward model $R(\boldsymbol{c},\mathbf{x}_{0})$. Simultaneously, the optimization imposes a regularization term that penalizes deviations from a reference distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})$. Formally, the objective function is formulated as follows:
$\begin{split}\max_{\mathbb{P}_{\theta}}\mathbb{E}_{\boldsymbol{c}\sim\mathcal{D}_{c},\mathbf{x}_{0}\sim\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})}\left[R(\boldsymbol{c},\mathbf{x}_{0})\right]\\ -\beta D_{\text{KL}}\left[\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})\parallel\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})\right]\,,\end{split}$ (8)
where $\beta$ is a hyper-parameter that governs the balance between the reward maximization and the regularization effect imposed by the KL-divergence. The regularization is crucial because it prevents the model from straying too far from the distribution where the reward model is reliable, while also preserving generation diversity and preventing mode collapse to a single high-reward sample. Replacing $R(\mathbf{x}_{0},\boldsymbol{c})$ with $R_{\text{NPO}}(\mathbf{x}_{0},\boldsymbol{c})$ yields the RLHF objective for Negative Preference Optimization (NPO):
$\begin{split}\max_{\mathbb{P}_{\theta}}\mathbb{E}_{\boldsymbol{c}\sim\mathcal{D}_{c},\mathbf{x}_{0}\sim\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})}\left[R_{\text{NPO}}(\boldsymbol{c},\mathbf{x}_{0})\right]\\ -\beta D_{\text{KL}}\left[\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})\parallel\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})\right]\,.\end{split}$ (9)
@@ 3.1 NPO via Self-Generated Data
Observing that a diffusion model can naturally produce undesirable outputs (e.g., disordered compositions, incoherent structures, blurry details), such low-quality samples inherently yield low reward scores. Consequently, exploiting self-generated data provides a straightforward path toward distribution-regularized negative preference optimization. Figure 2 summarizes the core idea and situates it among related methods. This approach is justified on two grounds: 1) Distribution preservation. Learning from self-generated data preserves the original distributional properties of the model. 2) Reward reduction. Self-generated samples exhibit low reward scores for several reasons: a) Hallucination. Generation of unrealistic objects, artifacts, or contextually inconsistent outputs, which result in low reward scores due to poor quality or implausibility of the generated content. b) Mode collapse. The model may fail to generate a diverse set of outputs, leading to a lack of coverage across the entire data distribution and, consequently, lower reward scores due to insufficient variety. c) Optimization. Imperfect optimization within the diffusion model itself can produce suboptimal results, such as blurry or noisy details, which leads to lower reward scores due to the misalignment between generated outputs and the desired target distribution. Indeed, we find it is helpful to intentionally corrupt the ODE sampling process to some extend, thus increasing the probability of reward reduction. We have more discussion in the Supplementary IV.1.
Figure 4: Truncated Diffusion Fine-Tuning. Baseline method requires full simulation of generation process, introducing huge amount of generation costs. Naive Truncation reduces the simulation costs but introduces distribution mismatch. Truncated Diffusion Fine-tuning reduces the simulation costs and maintains the target distribution.
@@
3.2 Truncated Diffusion Fine-tuning
@@
Our method: Truncated Diffusion Fine-tuning. According to the Tweedie’s formula, for the semi-implicit distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}^{ref}_{t}\mid\boldsymbol{c})$ induced by the reference diffusion model, the expected value of $\mathbf{x}_{0}$ given $\mathbf{x}_{t}^{ref}$, based on the conditional $\mathbb{P}(\mathbf{x}_{0}|\mathbf{x}_{t}^{ref})=\frac{\mathbb{P}(\mathbf{x}_{t}^{ref}|\mathbf{x}_{0})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}$ according to Bayes’ rule, is related to the score of $\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})$ efron2011tweedie ; zhou2024score . Specifically, the expectation is:
@@
[…]
@@
[…]
where $\mathbf{x}_{t^{\prime}}^{ref}$ is obtained by adding noise from real data following the forward process $\mathbb{P}(\mathbf{x}_{t}\mid\mathbf{x}_{0})$. When $t^{\prime}=T$ and $t=1$, we have $\mathbf{x}_{t^{\prime}}^{ref}=\mathbf{x}_{T}^{ref}$ and $\mathbf{x}_{t\rightarrow 0}^{ref}=\mathbf{x}_{1\rightarrow 0}^{ref}=\mathbf{x}_{0}^{ref}$. In this case, truncated diffusion fine-tuning essentially reduces to the baseline method. Thus, truncated diffusion fine-tuning can be viewed as a generalization of the baseline method, as it eliminates the need for fully solving the denoising process numerically, leading to significant improvements in training efficiency.
Figure 5: Visual Comparison on CogVideoX. Prompts are set as “A pretty lady at the beach and smiling, well drawn face” and “jake gyllenhaall drinking an ice cold beer, painting by egon schiele”. The results below or on the right are generated with self-npo, while the results above or on the left are without it.
@@
In the following, we theoretically show the reasonability of this operation from three aspects: 1) Distribution of $\mathbf{x}_{s}$. We show the the distribution of $\mathbf{x}_{s}$ is equivalent to the noise perturbed distribution of $\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})$ (See Theorem II.1). 2) Gradient equivalent optimization target. We show our optimization objective Equation 17 has an gradient equivalent learning objective, which is the expectation value of all $\mathbf{x}_{t\rightarrow 0}^{ref}$ (See Theorem II.2). 3) Equivalent optimization objective of diffusion models. We show the gradient equivalent learning objective as proved in 2) is equivalent to the learning objective of normal diffusion model (See Theorem II.3).
4 Experiments
4.1 Validation Setup
To better evaluate the performance of our method, we tested three baseline models: a) Stable Diffusion v1-5, a text-to-image generation model with approximately 860 million parameters, widely used for creating high-quality images from textual prompts in tasks like digital art and concept visualization; b) Stable Diffusion XL (SDXL), an advanced version with around 2.6 billion parameters, enabling higher-resolution image generation (e.g., 1024x1024) and improved compositional coherence; c) CogVideoX-5B, a video generation model with about 5 billion parameters, designed for tasks like video synthesis and frame interpolation, leveraging its larger capacity for temporal modeling. Therefore, our testing covers diffusion models of varying sizes and encompasses mainstream text-to-image and text-to-video tasks, ensuring a comprehensive and well-designed baseline.
4.2 Comparison
@@
Method PickScore HPSv2.1 ImageReward Aesthetic

### v2

# Self-NPO: Data-Free Diffusion Model Enhancement via Truncated Diffusion Fine-Tuning

@@ Abstract
Diffusion models have demonstrated remarkable success in various visual generation tasks, including image, video, and 3D content generation. Preference optimization (PO) is a prominent and growing area of research that aims to align these models with human preferences. While existing PO methods primarily concentrate on producing favorable outputs, they often overlook the significance of classifier-free guidance (CFG) in mitigating undesirable results. Diffusion-NPO addresses this gap by introducing negative preference optimization (NPO), training models to generate outputs opposite to human preferences and thereby steering them away from unfavorable outcomes through CFG. However, prior NPO approaches rely on costly and fragile procedures for obtaining explicit preference annotations (e.g., manual pairwise labeling or reward model training), limiting their practicality in domains where such data are scarce or difficult to acquire. In this work, we propose Self-NPO, specifically truncated diffusion fine-tuning, a data-free approach of negative preference optimization by directly learning from the model itself, eliminating the need for manual data labeling or reward model training. This data-free approach is highly efficient (less than 1% training cost of Diffusion-NPO) and achieves comparable performance to Diffusion-NPO in a data-free manner. We demonstrate that Self-NPO integrates seamlessly into widely used diffusion models, including SD1.5, SDXL, and CogVideoX, as well as models already optimized for human preferences, consistently enhancing both their generation quality and alignment with human preferences. Code is available at https://github.com/G-U-N/Diffusion-NPO.
Figure 1: Motivation of Self-NPO. Preference Optimization (PO) optimizes the preference aligned model by training on preference annotations. Negative Preference Optimization (NPO) optimizes the negative preference aligned model by training on switched preference annotations. Self-NPO optimizes the negative preference aligned model by truncated diffusion fine-tuning (TDFT) on truncated self-generated data.
Introduction
@@
Over the past few years, diffusion models (ho2020denoising; song2020denoising; song2020score) have made significant strides in various visual generation tasks, including image (karras2022elucidating; rombach2022high; wang2024rectified; podell2023sdxl; diffusionbeatgan; meng2021sdedit; teng2024dim; ma2024storynizor; li2024ecnet; sun2024journeydb), video (blattmann2023align; mao2024osv; shi2024motion; makeavideo; chen2024videocrafter2; bian2025gs; NEURIPS2022_f7f47a73; wang2024animatelcm; wang2025your; wang2024phased), and 3D generation (gao2024cat3d; poole2022dreamfusion; li2025connecting; yan2025consistent; sun2024genca; lai2025unleashing). However, diffusion models trained on massive unfiltered data (e.g., image-text pairs) often generate results that do not align well with human preferences. The growing interest in aligning these models for human preference-aligned generation has led to the development of preference optimization (PO) methods. In general, current PO methods can be categorized into four types: ($\mathrm{i}$) Differentiable Reward (DR) evaluates images from iterative denoising using a pre-trained reward model, refining diffusion models through gradient-based backpropagation to align with the reward model (xu2024imagereward; prabhudesai2024video; zhang2024large; wu2023hpsv1; wu2023human; wu2024deepreward; clark2023directly). ($\mathrm{ii}$) Reinforcement Learning (RL) models the denoising process as a Markov decision process, using techniques like PPO to optimize preferences by dynamically generating and evaluating images to maximize rewards (puterman2014markov; sutton2018reinforcement; schulman2017proximal). ($\mathrm{iii}$) Direct Preference Optimization (DPO) streamlines training by fine-tuning with paired preference datasets, eliminating separate reward models or online evaluation, often using high-quality data as positive examples and self-generated data as negatives (rafailov2024direct; wallace2024diffusion; yuan2024self; deng2024enhancing; gu2024diffusion; zhang2024seppo). ($\mathrm{iv}$) Negative Preference Optimization (Diffusion-NPO) trains models to avoid poor outputs by opposing human preferences, improving performance through classifier-free guidance (CFG) to steer away from undesired results (wang2025diffusionnpo; ho2022classifier; karras2024guiding; Shen_2024_CVPR; ahn2024self).
@@
However, we argue that a major limitation of existing methods is their reliance on data sources, ie, explicit preference annotations, such as the expensive manual collection of preference data pairs (kirstain2023pick; wu2023hpsv2) and the challenging and fragile process of reward model training (xu2024imagereward; wu2023hpsv2; schuhmann2022laion). This heavy dependency on preference annotations makes these approaches less practical, particularly in domains where obtaining or labeling preference data is expensive or infeasible.
@@
In this work, we introduce Self-NPO, specifically, truncated diffusion fine-tuning (TDFT), a novel approach for negative preference optimization (NPO) that can be done in a data-free manner. Specifically, our training objective is to apply negative preference optimization to a pre-trained model, making its predictions more aligned with the opposite direction of human preferences. This facilitates unconditional/negative-conditional outputs in classifier-free guidance, thus reducing the probability of generating results that do not align with human preferences.
@@
    To avoid altering the variance of the epsilon predictions, $\boldsymbol{\epsilon}_{\boldsymbol{\theta}_{neg}}(\mathbf{x}_{t},t,\boldsymbol{c^{\prime}})$ should maintain a certain level of correlation with $\boldsymbol{\epsilon}_{\boldsymbol{\theta}_{pos}}(\mathbf{x}_{t},t,\boldsymbol{c})$. For example, for two completely independent Gaussian noises, the above operation would result in a variance of $2\omega^{2}+2\omega+1$.
@@
    Efficient tuning without full diffusion. We do not need to fully execute iterative diffusion generation process to implement self-learning. We propose a novel tuning strategy, termed as truncated diffusion fine-tuning, which allows us to fine tune diffusion models on the partially excuted iterative diffusion generation results. This avoids the huge generation costs of preparing data for self-learning.
@@
We demonstrate that Self-NPO can be seamlessly integrated with existing models, such as SD1.5 (rombach2022high), SDXL (podell2023sdxl), and CogVideoX (yang2024cogvideox), as well as models previously fine-tuned through preference optimization, consistently improving their alignment with human preferences.
Preliminary
@@
Preliminary of CFG. CFG has become a necessary and important technique for improving generation quality and text alignment of diffusion models. For convenience, we focus our discussion on the general formal of diffusion models, i.e., $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\boldsymbol{\epsilon}$ (kingma2021variational). Suppose we learn a score estimator from a epsilon prediction neural network $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\mathbf{x}_{t},\boldsymbol{c},t)$, and we have $\nabla_{\mathbf{x}_{t}}\log\mathbb{P}_{\boldsymbol{\theta}}(\mathbf{x}_{t}|\boldsymbol{c};t)=-\frac{\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)}{\sigma_{t}}$. The sample prediction at timestep $t$ of the score estimator is formulated as
@@
$\hat{\mathbf{x}}_{0}=\frac{1}{\alpha_{t}}(\mathbf{x}_{t}+\sigma^{2}\nabla_{\mathbf{x}_{t}}\log\mathbb{P}_{\boldsymbol{\theta}}(\mathbf{x}_{t}|\boldsymbol{c};t))\,.$ (2)
@@
Applying the CFG is equivalent to add an additional score term, that is, we replace $\nabla_{\mathbf{x}_{t}}\log\mathbb{P}_{\boldsymbol{\theta}}(\mathbf{x}_{t}|\boldsymbol{c};t)$ in Eq. 2 with the following term,
[…]
@@
where $\omega$ is to control the strength of CFG, $\boldsymbol{c}$ and $\boldsymbol{c}^{\prime}$ are conditional and unconditional/negative-conditional inputs, respectively. It is apparent that the generation will be pushed to high probability region of $\mathbb{P}_{\boldsymbol{\theta}}(\mathbf{x}_{t}|\boldsymbol{c};t)$ and relatively low probability region of $\mathbb{P}_{\boldsymbol{\theta}}(\mathbf{x}_{t}|\boldsymbol{c}^{\prime};t)$. Write the above equation into the epsilon format, and then we have
$\boldsymbol{\epsilon}^{\omega}_{\boldsymbol{\theta}}=(\omega+1){\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\mathbf{x}_{t},\boldsymbol{c},t)}-\omega{\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\mathbf{x}_{t},\boldsymbol{c}^{\prime},t)}\,.$ (4)
@@
Diffusion-NPO (wang2025diffusionnpo). Diffusion models use classifier-free guidance (CFG) to improve generation quality by combining conditional and negative-conditional outputs to favor preferred results. Traditional preference optimization aligns outputs with human preferences but overlooks avoiding undesirable outputs. Negative Preference Optimization (NPO), introduced in Diffusion-NPO, trains an additional model to counteract human preferences, reducing undesired outputs via CFG.
@@
Figure 4: Truncated diffusion fine-tuning (TDFT). Baseline method requires full simulation of generation process, introducing huge amount of generation costs. Naive truncation reduces the simulation costs but introduces distribution mismatch. TDFT reduces the simulation costs and maintains the target distribution.
NPO via self-generated data
Reinforcement Learning with Human Feedback (RLHF) (rafailov2024direct; wallace2024diffusion; griffith2013policy; liu2025improving) aims to optimize a conditional distribution $\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})$, where $\boldsymbol{c}\sim\mathcal{D}_{c}$, in order to maximize the expectation value of the associated reward model $R(\mathbf{x}_{0},\boldsymbol{c})$. Simultaneously, the optimization imposes a regularization term that penalizes deviations from a reference distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})$. Formally, the objective function is formulated as follows:
[…]
where $\beta$ is a hyper-parameter that governs the balance between the reward maximization and the regularization effect imposed by the KL-divergence. The regularization is crucial because it prevents the model from straying too far from the distribution where the reward model is reliable, while also preserving generation diversity and preventing mode collapse to a single high-reward sample. Replacing $R(\mathbf{x}_{0},\boldsymbol{c})$ with $R_{\text{NPO}}(\mathbf{x}_{0},\boldsymbol{c})$ yields the learning objective for Negative Preference Optimization (NPO):
$\begin{split}\max_{\mathbb{P}_{\theta}}\mathbb{E}_{\boldsymbol{c}\sim\mathcal{D}_{c},\mathbf{x}_{0}\sim\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})}\left[R_{\text{NPO}}(\mathbf{x}_{0},\boldsymbol{c})\right]\\ -\beta D_{\text{KL}}\left[\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})\parallel\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})\right]\,.\end{split}$ (9)
@@
Observing that a diffusion model can naturally produce undesirable outputs (e.g., disordered compositions, incoherent structures, blurry details), such low-quality samples inherently yield low reward scores. Consequently, exploiting self-generated data provides a straightforward path toward distribution-regularized negative preference optimization. Fig. 1 summarizes the core idea and situates it among related methods. This approach is justified on two grounds: 1) Distribution preservation. Learning from self-generated data preserves the original distributional properties of the model. 2) Reward reduction. Self-generated samples exhibit low reward scores for several reasons: a) Hallucination. Generation of unrealistic objects, artifacts, or contextually inconsistent outputs, which result in low reward scores due to poor quality or implausibility of the generated content. b) Mode collapse. The model may fail to generate a diverse set of outputs, leading to a lack of coverage across the entire data distribution and, consequently, lower reward scores due to insufficient variety. c) Optimization. Imperfect optimization within the diffusion model itself can produce suboptimal results, such as blurry or noisy details, which leads to lower reward scores due to the misalignment between generated outputs and the desired target distribution. Indeed, we find it is helpful to intentionally corrupt the ODE sampling process to some extent, thus increasing the probability of reward reduction. We have more discussion and empirical evidence in the supplementary material.
@@
Truncated diffusion fine-tuning
@@
Our method: Truncated Diffusion Fine-tuning. According to the Tweedie’s formula, for the semi-implicit distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}^{ref}_{t}\mid\boldsymbol{c})$ induced by the reference diffusion model, the expected value of $\mathbf{x}_{0}$ given $\mathbf{x}_{t}^{ref}$, based on the conditional $\mathbb{P}(\mathbf{x}_{0}|\mathbf{x}_{t}^{ref})=\frac{\mathbb{P}(\mathbf{x}_{t}^{ref}|\mathbf{x}_{0})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}$ according to Bayes’ rule, is related to the score of $\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})$ (efron2011tweedie; zhou2024score). Specifically, the expectation is:
@@
$\mathbf{x}^{ref}_{T}\rightarrow\mathbf{x}^{ref}_{T-1}\rightarrow\dots\mathbf{x}^{ref}_{t}\xrightarrow{\text{Tweedie's formula}}\mathbf{x}_{t\rightarrow 0}^{ref}\,$, (13)
@@
$\mathbf{x}^{ref}_{t^{\prime}}\rightarrow\mathbf{x}^{ref}_{t^{\prime}-1}\rightarrow\dots\mathbf{x}^{ref}_{t}\xrightarrow{\text{Tweedie's formula}}\mathbf{x}_{t\rightarrow 0}^{ref}\,$, (14)
where $\mathbf{x}_{t^{\prime}}^{ref}$ is obtained by adding noise from real data following the forward process $\mathbb{P}(\mathbf{x}_{t}\mid\mathbf{x}_{0})$. When $t^{\prime}=T$ and $t=1$, we have $\mathbf{x}_{t^{\prime}}^{ref}=\mathbf{x}_{T}^{ref}$ and $\mathbf{x}_{t\rightarrow 0}^{ref}=\mathbf{x}_{1\rightarrow 0}^{ref}=\mathbf{x}_{0}^{ref}\$. In this case, truncated diffusion fine-tuning essentially reduces to the baseline method. Thus, truncated diffusion fine-tuning can be viewed as a generalization of the baseline method, as it eliminates the need for fully solving the denoising process numerically, leading to significant improvements in training efficiency.
@@
Theoretical grounding
We theoretically show the reasonability of proposed TDFT from three aspects:
  1. 1.
    Distribution of $\mathbf{x}_{s}$. We show that the distribution of $\mathbf{x}_{s}$ is equivalent to the noise-perturbed distribution of $\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{\text{ref}}\mid\boldsymbol{c})$ (See Thm. 1).
  2. 2.
    Gradient equivalent optimization target. We show that our optimization objective, Eq. (17), has a gradient-equivalent learning objective, which is the expectation value of all $\mathbf{x}_{t\rightarrow 0}^{\text{ref}}$ (See Thm. 2).
  3. 3.
    Equivalent optimization objective of diffusion models. We show that the gradient-equivalent learning objective, as proved in 2), is equivalent to the learning objective of a standard diffusion model (See Thm. 3).
Proof Sketch: Due to the limited space, we encourage readers to review the detailed proofs in Thm. 1, Thm. 2, and Thm. 3. These proofs rigorously establish the equivalence of distributions, optimization objectives, and their connections to diffusion models, forming the foundation for understanding the validity and effectiveness of our proposed approach.
Experiments
Validation setup
To better evaluate the performance of our method, we tested three baseline models: a) Stable Diffusion v1-5, a text-to-image generation model with approximately 860 million parameters, widely used for creating high-quality images from textual prompts in tasks like digital art and concept visualization; b) Stable Diffusion XL (SDXL), an advanced version with around 2.6 billion parameters, enabling higher-resolution image generation (e.g., 1024x1024) and improved compositional coherence; c) CogVideoX-5B, a video generation model with about 5 billion parameters, designed for tasks like video synthesis and frame interpolation, leveraging its capacity for temporal modeling. Therefore, our testing covers diffusion models of varying sizes and encompasses mainstream text-to-image and text-to-video tasks, ensuring a comprehensive and well-designed baseline.
Comparison
@@
Method P.S. HPS. I.R. AES.